There are 7 girls in a bus. Each girl has 7 backpacks. In each backpack, there are 7 big cats. For every big cat, there are 7 little cats. The bus driver is not in the bus at this time.
Question:
How many legs are there in the bus?

Answer by checkley77(12569)

7 girls=14 legs
7girls*7backpacks*7big cats=7*7*7*4=343*4=1,372 big cat legs.
343*7*4=9,604 little cat legs
14+1,372+9,604=10.990 total number of legs.